OpenStudy (anonymous):

Solve 3x2 = –30x + 24.

OpenStudy (anonymous):

x^2

OpenStudy (anonymous):

\[-5 \pm \sqrt{33}\]

OpenStudy (anonymous):

also, Solve 3x2 + 4x = 2.

OpenStudy (anonymous):

\[1/3 (-2 \pm \sqrt{10})\]